OpenStudy (anonymous):

Last year Lien earned $1,300 more than her husband.Together they earned $59,600.how much did each of them earn?

OpenStudy (anonymous):

Sandra, can you help me with my fractions homework? It is 5th grade math. (fractions)

OpenStudy (sandra):

l = h + 1300 l + h = 59600 substitute the first into the second: h + 1300 + h = 59600 2h + 1300 = 59600 2h = 58300 h = 29650 now substitute back into the first equation to get l l = 1300 + h l = 1300 + 29650 l = 30,950

OpenStudy (sandra):

hey nya, can you ask a question on the left hand side?

OpenStudy (sandra):

(to start a new conversation)

OpenStudy (anonymous):

Yes I can!

OpenStudy (anonymous):

Thanks!

OpenStudy (sandra):

np, I'll jump in